HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

ISOTOPIC SUBSTITUTION IN POLYATOMIC MOLECULES: COMPARATIVE LINE POSITION AND LINE STRENGTH ANALYSIS OF

THE ν_{2}/ν_{4} DYAD OF^{12}CD_{4} AND^{13}CD_{4}

\underline{O. N. ULENIKOV} , \underline{O. V. GROMOVA} , \underline{E. S. BEKHTEREVA} ,

\underline{N. I. NIKOLAEVA} , National Research Tomsk Polytechnic University, 30, av. Lenina, 634050 Tomsk, Russia; \underline{C. SYDOW}, \underline{C. MAUL} and \underline{S. BAUERECKER} ,

Institut für Physikalische und Theoretische Chemie, Technische Universität Braunschweig, D - 38106, Braunschweig, Germany

A highly accurate ro–vibrational analysis of FTIR spectra (line positions of^{13}CD_{4} and line strengths of both the ^{12}CD_{4} and ^{13}CD_{4} species) is presented. The high– resolution infrared spectra of both molecules were measured with a Bruker IFS125 HR Fourier transform infrared spectrometer at an optical resolution of 0.003 cm^{−1} and analyzed in the regions of 800–1400 cm^{−1} where the ν_{2}/ν_{4} dyad is located. The number of 901 transitions with Jmax = 23 were assigned to the ν_{4} and ν_{2} bands of ^{13}CD_{4}. The weighted fit of experimental line positions was made using the Hamiltonian model which takes into account the resonance interactions between the (0001, F_{2}) and (0100, E) vibrational states. As a result, set of 18 fitted parameters of the (0001, F_{2})/(0100, E) vibrational states of ^{13}CD_{4} was determined which reproduce the initial 901 experimental ro–vibrational line positions with the drms = 2.59 × 10^{−4} cm^{−1}, which is close to the experimental uncertainty of the recorded spectra. The analysis of 1557 experimental lines of the dyad of ^{12}CD_{4} and 131 lines of the dyad of ^{13}CD_{4} was fulfilled with the Hartmann–Tran profile to simulate the measured line shape and to determine experimental line intensities. Sets of 6/1 varied effective dipole moment parameters of ^{12}CD_{4}/^{13}CD_{4} are determined which reproduce the initial 1557/131 line strengths with the drms = 4.80% and 4.21%.

The study was financially supported by Volkswagen Foundation.

p-number: p016  Submitted on Sat, 27 May 2023 15:03:00 +0200